OpenStudy (anonymous):

Simplify 3^-9 / 3^-12

OpenStudy (anonymous):

answers: 3^3 3^21 1/3^21 1/3^3

OpenStudy (asnaseer):

do you know the rule:\[x^a/x^b=x^{a-b}\]

OpenStudy (anonymous):

nope sure dont

OpenStudy (asnaseer):

what have you been taught about indexes?

OpenStudy (anonymous):

nothing at all that i can recall

OpenStudy (anonymous):

no pun intended

OpenStudy (asnaseer):

so how do they expect you to solve this? have you been told to use any specific method?

OpenStudy (anonymous):

i think we have i just dont remember

OpenStudy (asnaseer):

well at the very least you need to know the rules of index multiplication and division.

OpenStudy (anonymous):

prolly do just dont remember what they are

OpenStudy (asnaseer):

e.g. the rule I just showed you can be proved with following example:\[\frac{x^5}{x^3}=\frac{x*x*x*x*x}{x*x*x}=x*x=x^2\]

OpenStudy (anonymous):

oooo thtz right i remember now thx

OpenStudy (asnaseer):

so in your case,\[3^{-9}/3^{-12}=3^{-9-(-12)}=3^{-9+12}=3^3\]

OpenStudy (anonymous):

yup i got it thanks alot

OpenStudy (asnaseer):

yw